# RegMixMatch: Optimizing Mixup Utilization in Semi-Supervised Learning

### v1

@@ Introduction
Fully-Supervised 4.62$\pm$0.05 19.30$\pm$0.09 2.13$\pm$0.01 -
@@
In addition to utilizing low-confidence samples, a key distinction of RegMixMatch from other Mixup-based SSL methods is its retention of clean samples within SRM during training. To verify the rationale behind this design, we compare RegMixMatch with two scenarios: one where only clean samples are retained during training (w/o mixed samples) and another where only mixed samples are used (w/o clean samples). To further validate the rationale and effectiveness of CAM in leveraging low-confidence data, we also present results for three additional configurations: removing CAM entirely (w/o CAM), replacing CAM with Mixup (CAM $\rightarrow$ Mixup) and adapting the CAM to incorporate information from the top-3 predicted classes (Top-2 $\rightarrow$ Top-3). The results for these ablation studies on STL-10 with 40 labels are shown in Table 3.
@@
Table 3 highlights several key findings: 1) Training with only clean samples or only mixed samples yields suboptimal results compared to using both. RegMixMatch leverages SRM not only to enhance generalization performance but also to mitigate the negative effects of high-entropy behavior in SSL, as shown in Figure 4(a). 2) CAM improves both classification accuracy and learning efficiency, as evidenced by the higher accuracy at the same iterations in Figure 4(b). This suggests that CAM facilitates the safer utilization of low-confidence samples, enriching the training process with more informative data. 3) Replacing CAM with Mixup or incorporating features from more specific classes (i.e., Top-3) results in performance degradation. This underscores the importance of specifically limiting the classes and quantities of mixed samples for low-confidence data and demonstrates the rationale of CAM.
@@
In addition to retaining clean samples, RegMixMatch distinguishes itself from previous Mixup-based SSL methods by generating more robust mixed samples to enhance model generalization. The parameter $\alpha$ controls the mixing intensity, typically set to a value less than 1.0 in prior work. However, as shown in Figures 5(b) and 5(c), the optimal values for $\alpha_{h}$ and $\alpha_{l}$ in RegMixMatch are 1.0 and 16.0, respectively. A higher $\alpha$ indicates stronger augmentation, improving generalization performance. It is noteworthy that $\alpha_{l}\gg\alpha_{h}$. As observed with Mixup, a greater mixing intensity for noisy data enhances the model’s robustness to noise. Since the supervisory information for low-confidence samples is more prone to noise, a higher $\alpha_{l}$ generates more intermediate-state samples, increasing the quantity and variability of the training data, and thereby reducing the risk of overfitting to specific noisy samples.
@@ Acknowledgments
This work was supported by the Fundamental Research Funds for the Central Universities (No.2022JBMC011).

### v2

@@ Introduction
VAT (Miyato et al. 2018) 79.81 74.66 41.03 10.51 85.20 46.84 32.14 74.75 4.33 4.11 74.74 37.95
Mean Teacher (Tarvainen and Valpola 2017) 76.37 70.09 37.46 8.10 81.11 45.17 31.75 36.09 3.45 3.27 71.72 33.90
MixMatch (Berthelot et al. 2019b) 65.76 36.19 13.63 6.66 67.59 39.76 27.78 30.60 4.56 3.69 54.93 21.70
ReMixMatch (Berthelot et al. 2019a) 20.77 9.88 6.30 4.84 42.75 26.03 20.02 24.04 6.36 5.16 32.12 6.74
UDA (Xie et al. 2020) 34.53 10.62 5.16 4.29 46.39 27.73 22.49 5.12 1.92 1.89 37.42 6.64
FixMatch (Sohn et al. 2020) 24.79 7.47 4.86 4.21 46.42 28.03 22.20 3.81 2.02 1.96 35.97 6.25
Dash (Xu et al. 2021) 27.28 8.93 5.16 4.36 44.82 27.15 21.88 2.19 2.04 1.97 34.52 6.39
MPL (Pham et al. 2021) 23.55 6.62 5.76 4.55 46.26 27.71 21.74 9.33 2.29 2.28 35.76 6.66
FlexMatch (Zhang et al. 2021) 13.85 4.97 4.98 4.19 39.94 26.49 21.90 8.19 6.59 6.72 29.15 5.77
SoftMatch (Chen et al. 2023) - 4.91 4.82 4.04 37.10 26.66 22.03 2.33 - 2.01 21.42 5.73
FreeMatch (Wang et al. 2023) 8.07 4.90 4.88 4.10 37.98 26.47 21.68 1.97 1.97 1.96 15.56 5.63
SequenceMatch (Nguyen 2024) - 4.80 4.75 4.15 37.86 25.99 20.10 1.96 1.89 1.79 15.45 5.56
FlatMatch (Huang et al. 2023) 15.23 5.58 4.22 3.61 38.76 25.38 19.01 2.46 1.43 1.41 16.20 4.82
RegMixMatch 4.35 4.24 4.21 3.38 35.27 23.78 19.41 1.81 1.77 1.79 11.74 4.66
Fully-Supervised 4.62 19.30 2.13 -
Table 1: Error rates on CIFAR10/100, SVHN, and STL10 datasets. The fully-supervised results of STL10 are unavailable since we do not have label information for its unlabeled data. The best results are highlighted with Bold and the second-best results are highlighted with underline. We ran each task three times, and the results with standard deviation are presented in the appendix.
@@
In addition to utilizing low-confidence samples, a key distinction of RegMixMatch from other Mixup-based SSL methods is its retention of clean samples within SRM during training. To verify the rationale behind this design, we compare RegMixMatch with two scenarios: one where only clean samples are retained during training (w/o mixed samples) and another where only mixed samples are used (w/o clean samples). To further validate the rationale and effectiveness of CAM in leveraging low-confidence data, we also present results for two additional configurations: removing CAM entirely (w/o CAM), and replacing CAM with Mixup (CAM $\rightarrow$ Mixup). The results for these ablation studies on STL-10 with 40 labels are shown in Table 3.
@@
Table 3 highlights several key findings: 1) Training with only clean samples or only mixed samples yields suboptimal results compared to using both. RegMixMatch leverages SRM not only to enhance generalization performance but also to mitigate the negative effects of high-entropy behavior in SSL, as shown in Figure 4(a). 2) CAM improves both classification accuracy and learning efficiency, as evidenced by the higher accuracy at the same iterations in Figure 4(b). This suggests that CAM facilitates the safer utilization of low-confidence samples, enriching the training process with more informative data. 3) Replacing CAM with Mixup results in performance degradation.
@@
In addition to retaining clean samples, RegMixMatch distinguishes itself from previous Mixup-based SSL methods by generating more robust mixed samples to enhance generalization. The parameter $\alpha$ typically set to a value less than 1.0 in prior work. However, as shown in Figures 5(b) and 5(c), the optimal values for $\alpha_{h}$ and $\alpha_{l}$ in RegMixMatch are 1.0 and 16.0, respectively. A higher $\alpha$ indicates stronger augmentation, improving generalization performance. It is noteworthy that $\alpha_{l}\gg\alpha_{h}$. As observed with Mixup, a greater mixing intensity enhances the model’s robustness to noise. Since the supervisory information for low-confidence samples is more prone to noise, a higher $\alpha_{l}$ generates more intermediate-state samples, increasing the quantity and variability of the training data, and thereby reducing the risk of overfitting to specific noisy samples.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62402031) and the Beijing Nova Program (No. 20240484620).
@@ Appendix B Additional Experiment Results
Dataset CIFAR10 CIFAR100 SVHN STL10
# Label 10 40 250 4000 400 2500 10000 40 250 1000 40 1000
Fully-Supervised 4.62$\pm$0.05 19.30$\pm$0.09 2.13$\pm$0.01 -